OpenStudy (anonymous):

Use Binomial Theorem to expand the binomial. (x + 2y^2)^3

OpenStudy (anonymous):

(a+b)^3 = a^3 + 3a^2b + 3ab^2 + b^3 where the coefficients of the terms come from Pascal's Triangle or Combinations. In this problem a = x and b = 2y^2.

OpenStudy (isaiah.feynman):

\[(a+b)^{3}=a^{3}+3a^{2}b+3ab^{2}+b^{3} \]

OpenStudy (anonymous):

Thank you everyone!